The following stem-and-leaf chart reports the number of movies rented per day at Video Connection:

12 1 2 3 13 14 6 889 15 3 6 6 16 35 17 2 4568 18 19 1 345 6 20 34679 21 2 2 3 9 22 78 9 23 00 179 24 8 25 1 26 27 1 2 3 2 6 8 1 2 3 3 4 5 667

Questions:
Answer #1

Here' the answer to the problem in detail. please don't hesitate to give a "thumbs up " for the answer in case you're satisfied with the answer.

a.

This will be total obs in the stem-leaf diagram:
=(3+3+4+3++2+5+3+5+6+4+3+5+1+2+0+9)
= 58
b. In the last class there are 9 observations

c. Smallest and largest values in the entire set is 121 and 277 respectively

d. In the fourth row the values are : 153, 156, 156

e. question not valid

f. Anything less than 160 movies is : 3+3+4+3 = 13 days

g. Anything >= 220 days is 20 days. Answer is 20 days

h. As there are 59 values, mid of 29th and 30th is 200+203/2 = 201.5. This is the answer.

i. between 240 and 280 there are 2+0+9 = 11 movies